# Continual Segmentation with Disentangled Objectness Learning and Class Recognition

### v2

@@ 2 Related Work
Query-based Image Segmentation. Query-based segmenters [9, 10, 55, 21] have unified image segmentation via mask classification and solve semantic, instance, and panoptic segmentation using the same framework. However, the mask classification paradigm was originated from query-based detectors, such as DETR [2] and its variants [66, 24, 24, 58], where we train the model to predict proposals and class labels with learnable queries. MaskFormer [9] was the first to introduce mask classification in semantic segmentation, which was dominated by per-pixel classification paradigm via convolutional nets [8, 52, 46, 62] or Transformers [62, 42, 53]. Following MaskFormer, Mask2Former [10] introduced multiple improvements, such as masked attention for mask refinement and using multi-scale features for better small object performance. To speed up the convergence, $k$MaX-DeepLab [55] generated queries from clustered image features. Recently, OneFormer [21] provided a generalist solution by joint training of all segmentation tasks. However, all of these segmenters learn to predict to mask and class simultaneously and are not suitable for decoupled continual segmentation.
Continual Segmentation. Learning new knowledge by fintuning the model is a native way, but resulting catastrophic forgetting of old knowledge [16, 31, 39]. Continual learning aims to alleviate the forgetting issue by balancing the learning of new knowledge (plasticity) and the preserving of the old one (stability). While the research focus of continual learning is on classification [22, 25, 38, 1, 11, 6, 50, 41], continual segmentation started to thrive recently [3, 12, 32, 7, 36, 57, 59, 5, 40, 51, 65, 60]. As pointed out in MiB [3], the background shift in continual segmentation aggravates the forgetting issue. Cermelli et al. [3] proposed unbiased distillation and classification to mitigate the background shift. PLOP [12] circumvented the background shift issue by feature distillation and pseudo-labeling. Based on PLOP, REMINDER [36] introduced class-weighted knowledge distillation to discriminate the new and old classes. With the same motivation, Incrementer [40] designed a class deconfusion strategy and achieved competitive performance via a Transformer-based segmenter. Recently, RL-Replay [65] set a strong replay-based baseline via reinforcement learning. Some works [4, 20] started to focus on more realistic scenarios by integrating weak supervision [61, 56] with continual learning. The methods above follow the per-pixel paradigm. Alhough CoMFormer [5] was the first to present mask classification in continual segmentation, it applied a standard framework with distillation and pseudo-labeling and failed to leverage the benefit of objectness. In contrast, CoMasTRe exploits objectness and decouples continual segmentation into forgetting-resistant objectness learning and class recognition.
@@ 4.2 Quantitative Evaluation
PASCAL VOC 2012. We compare our method with previous state-of-the-art methods on the 3 settings on PASCAL VOC: 19-1, 15-5, and 15-1 in Tab. 1. In general, CoMasTRe significantly outperforms query-based CoMFormer [5] and surpasses per-pixel baselines on incremented classes by a large margin. On 19-1, CoMasTRe shows off its strong plasticity in learning new classes, with 32.16 percent points (p.p) improvement while maintaining the knowledge of old classes. On 15-5, our model is on par with RCIL [57] on new classes while exceptionally great at maintaining old class performance, even surpassing the joint upper bound of per-pixel baselines by 0.85 p.p. 15-1 is much more difficult and involves 6 learning steps. However, CoMasTRe still beats RCIL with 4.14 p.p uplift on all classes. Overall, we strikes a balance between stability and plasticity.
ADE20K. As shown in Tab. 2, we report the results on ADE20K in 100-50, 100-10, and 100-5. In general, CoMasTRe outperforms previous state-of-the-art CoMFormer [5] in each setting by 0.8 p.p, 2.11 p.p, and 1.65 p.p on all classes, respectively. In 100-50, our method shows competitive performance with CoMFormer on incremented classes but with 0.8 p.p gain on all classes thanks to its higher upper bound. For longer learning processes, CoMasTRe reaches a comparable performance with REMINDER [36] on new classes but significantly boosts base class mIoU (3.36 p.p in 100-10 and 4.76 p.p in 100-5). Compared with CoMFormer, our CoMasTRe also learns better new classes with 2.82 p.p improvement in 100-10 and 2.23 p.p in 100-5 while preserving better old class knowledge, with 1.72 p.p and 1.32 p.p gains, respectively.
@@ 4.3 Ablation Studies
Joint training results. To resolve scalability concerns on mask-only matching, we demonstrate the joint performance of CoMasTRe, i.e., single-shot training on all classes. We compare the segmentation performance in mIoU on PASCAL VOC and ADE20k using ResNet50 (R50) and ResNet101 (R101) backbones in Tab. 3. The results show that CoMasTRe achieves comparable performance with Mask2Former and outperforms the previous continual model CoMFormer by $\sim$1.2 p.p on ADE20K.

### v3

@@ 2 Related Work
Query-based Image Segmentation. Query-based segmenters [9, 10, 55, 21] have unified image segmentation via mask classification and solve semantic, instance, and panoptic segmentation using the same framework. However, the mask classification paradigm was originated from query-based detectors, such as DETR [2] and its variants [66, 24, 59], where we train the model to predict proposals and class labels with learnable queries. MaskFormer [9] was the first to introduce mask classification in semantic segmentation, which was dominated by per-pixel classification paradigm via convolutional nets [8, 52, 46] or Transformers [42, 53]. Following MaskFormer, Mask2Former [10] introduced multiple improvements, such as masked attention for mask refinement and using multi-scale features for better small object performance. To speed up the convergence, $k$MaX-DeepLab [55] generated queries from clustered image features. Recently, OneFormer [21] provided a generalist solution by joint training of all segmentation tasks. However, all of these segmenters learn to predict to mask and class simultaneously and are not suitable for decoupled continual segmentation.
Continual Segmentation. Learning new knowledge by fintuning the model is a native way, but resulting catastrophic forgetting of old knowledge [16, 31, 39]. Continual learning aims to alleviate the forgetting issue by balancing the learning of new knowledge (plasticity) and the preserving of the old one (stability). While the research focus of continual learning is on classification [22, 25, 38, 1, 11, 6, 50, 41, 58], continual segmentation started to thrive recently [3, 12, 32, 7, 36, 57, 60, 5, 40, 51, 65, 61]. As pointed out in MiB [3], the background shift in continual segmentation aggravates the forgetting issue. Cermelli et al. [3] proposed unbiased distillation and classification to mitigate the background shift. PLOP [12] circumvented the background shift issue by feature distillation and pseudo-labeling. Based on PLOP, REMINDER [36] introduced class-weighted knowledge distillation to discriminate the new and old classes. With the same motivation, Incrementer [40] designed a class deconfusion strategy and achieved competitive performance via a Transformer-based segmenter. Recently, RL-Replay [65] set a strong replay-based baseline via reinforcement learning. Some works [4, 20] started to focus on more realistic scenarios by integrating weak supervision [62, 56] with continual learning. The methods above follow the per-pixel paradigm. Alhough CoMFormer [5] was the first to present mask classification in continual segmentation, it applied a standard framework with distillation and pseudo-labeling and failed to leverage the benefit of objectness. In contrast, CoMasTRe exploits objectness and decouples continual segmentation into forgetting-resistant objectness learning and class recognition.
@@ 4.2 Quantitative Evaluation
PASCAL VOC 2012. We compare our method with previous state-of-the-art methods on the 3 settings on PASCAL VOC: 19-1, 15-5, and 15-1 in LABEL:tab:voc. In general, CoMasTRe significantly outperforms query-based CoMFormer [5] and surpasses per-pixel baselines on incremented classes by a large margin. On 19-1, CoMasTRe shows off its strong plasticity in learning new classes, with 32.16 percent points (p.p) improvement while maintaining the knowledge of old classes. On 15-5, our model is on par with RCIL [57] on new classes while exceptionally great at maintaining old class performance, even surpassing the joint upper bound of per-pixel baselines by 0.85 p.p. 15-1 is much more difficult and involves 6 learning steps. However, CoMasTRe still beats RCIL with 4.14 p.p uplift on all classes. Overall, we strikes a balance between stability and plasticity.
ADE20K. As shown in LABEL:tab:ade, we report the results on ADE20K in 100-50, 100-10, and 100-5. In general, CoMasTRe outperforms previous state-of-the-art CoMFormer [5] in each setting by 0.8 p.p, 2.11 p.p, and 1.65 p.p on all classes, respectively. In 100-50, our method shows competitive performance with CoMFormer on incremented classes but with 0.8 p.p gain on all classes thanks to its higher upper bound. For longer learning processes, CoMasTRe reaches a comparable performance with REMINDER [36] on new classes but significantly boosts base class mIoU (3.36 p.p in 100-10 and 4.76 p.p in 100-5). Compared with CoMFormer, our CoMasTRe also learns better new classes with 2.82 p.p improvement in 100-10 and 2.23 p.p in 100-5 while preserving better old class knowledge, with 1.72 p.p and 1.32 p.p gains, respectively.
@@ 4.3 Ablation Studies
Joint training results. To resolve scalability concerns on mask-only matching, we demonstrate the joint performance of CoMasTRe, i.e., single-shot training on all classes. We compare the segmentation performance in mIoU on PASCAL VOC and ADE20k using ResNet50 (R50) and ResNet101 (R101) backbones in LABEL:tab:joint-bench. The results show that CoMasTRe achieves comparable performance with Mask2Former and outperforms the previous continual model CoMFormer by $\sim$1.2 p.p on ADE20K.